OpenStudy (unklerhaukus):

You are dealt two playing cards from a standard deck (without replacement). What is the probability that one of the cards is the 4 of diamonds?

Parth (parthkohli):

I \(\heartsuit \) probability.

OpenStudy (unklerhaukus):

\[\boxed{\color{red}{4\diamond}}\]

Parth (parthkohli):

P(the first card is four of diamonds) = 1/52. P(the second card is four of diamonds) = 51/52 * 1/51 Add both.

OpenStudy (unklerhaukus):

I get 2/52

Parth (parthkohli):

Good.

OpenStudy (unklerhaukus):



Parth (parthkohli):

yes, I think you can think of it like this. Is 1/26 right though?

OpenStudy (unklerhaukus):

This was my first thought,

OpenStudy (unklerhaukus):

1-51/52*50/51 = 1 - 50/52 = 2/52

OpenStudy (unklerhaukus):

hmm ok, i guess it really was that easy

OpenStudy (unklerhaukus):

i've been studying too much QM, no quantum interference here

OpenStudy (unklerhaukus):

If instead there is replacement, the probability is greater right? 1/52 + 51/52 * 1/52 + 1/52*1/52 = 2/52 + 1/52^2 > 2/52

ganeshie8 (ganeshie8):

w/ replacement shouldn't it be simply http://www.wolframalpha.com/input/?i=1+-+%2851%2F52%29%5E2

ganeshie8 (ganeshie8):

P("at least one time 4diamond") = 1 - P("no time 4diamond")

OpenStudy (unklerhaukus):

but 1/26 = 0.03846 > 0.03809

ganeshie8 (ganeshie8):

Now that looks very interesting...

OpenStudy (unklerhaukus):

what formulae are we using here?

Parth (parthkohli):

The probability of getting it with replacement should obviously be lower, no?

OpenStudy (unklerhaukus):

hmm, it's not so obvious

OpenStudy (unklerhaukus):

Why is is less?

Parth (parthkohli):

One more card in the deck makes it harder to get the intended one.

ganeshie8 (ganeshie8):

what if the replacement card itself is the intended one ?

Parth (parthkohli):

but that is only one case compared to the fifty-one others.

OpenStudy (ikram002p):

1/52+1/51

OpenStudy (unklerhaukus):

The chance of getting 4 of diamonds twice, is relatively insignificant.

ganeshie8 (ganeshie8):

`If instead there is replacement, the probability is greater right?` 1/52*51/52 + 51/52*1/52 + 1/52*1/52 = (2*51+1)/52^2

OpenStudy (unklerhaukus):

ah @ganeshie8, that's it and the previous case was 1/52 + 1/52 = 2*52/52^2 comparing numerators 2*52 > 2*51+1

Parth (parthkohli):

oh so was I wrong

OpenStudy (unklerhaukus):

I understand it better now, thanks to all

OpenStudy (unklerhaukus):

you were not wrong at all @ParthKohli

OpenStudy (unklerhaukus):

1/26 = 2*52/52^2

Parth (parthkohli):

yes, got it